Using a table of critical​ t-values of the t​ distribution, find the range of values for the​ P-value for a​ two-tailed test with

nequals=15

and test statistic t equals 1.976

What is the range of values for the​ P-value?

A.

0.01 less than Upper P dash value less than 0.020.01<P-value<0.02

B.

0.10 less than Upper P dash value less than 0.200.10<P-value<0.20

C.

0.025 less than Upper P dash value less than 0.050.025<P-value<0.05

D.

0.05 less than Upper P dash value less than 0.10
Answer #1

Solution,

Given that,

n = 15

degrees of freedom = n - 1 = 15 - 1 = 14

t = 1.976

P( t > 1.976 ) = 1 - P( t < 1.976 ) = 1 - 0.9659 = 0.0341

This is two tailed test,

P-value = 2 * P( t > 1.976 )

P-value = 2 * 0.0341

P-value = 0.0682

correct option is = D

0.05 < p-value < 0.10